The car is moving at a speed of 50 km / h. At one point the driver starts to brake abruptly and the car stops after 35 m from the moment the driver applied the brakes. Determine the car acceleration intensity during the braking process.


1
Expert's answer
2021-03-21T11:24:20-0400

We can find the car acceleration as follows:


"v^2=v_0^2+2ad,""a=\dfrac{v^2-v_0^2}{2d},""a=\dfrac{0-(13.88\ \dfrac{m}{s})^2}{2\cdot35\ m}=-2.75\ \dfrac{m}{s^2}."

The sign minus means that the car decelerates.